Ticket Number: HEA-NOC/20070615-3 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20070615 16:45 UTC+1 Problem Start: 20070627 09:00 UTC+1
Ticket Closed: 20070628 20:40 UTC+1 Problem End: 20070627 14:00 UTC+1
Adding new stack to Co-lo switch and upgrading IOS on blanch-colo1.
Co-location customers connected will experience an outage while the switch
is rebooted, but downtime will be kept to a minimum.
20070615 10:15 OMG Opening ticket to track work. The new version of IOS
will be installed on Monday 18th afternoon in advance of the maintenance
on Tuesday morning when the router will be reloaded. Scheduling 3 hours
for the work, but don't expect downtime greater than 5 minutes.
20070618 17:40 OMG Work has been postponed, as switch has not arrived in
time to Blanchardstown. Will reschedule work once a switch is delivered.
20070621 15:04 OMG Work has been rescheduled to Wednesday morning 27th
June. Updating co-location clients.
20070627 10:17 OMG New stack is racked. Beginning configuration and IOS
upgrade of switch.
20070627 13:03 OMG Maintenance window over-ran by 1 hour. Some further
configuration of the newly added switch stack is required. The other
switch stacks are functioning correctly - no further downtime is needed,
however the window has been extended until 14:00.
20070627 13:32 OMG Work complete. Will keep ticket open for monitoring.
20070628 15:36 OMG Switch is stable. Closing ticket.
This ticket can be monitored at http://www.hea.net/tickets/20070615-3
HEAnet Limited. Registered in Ireland, No. 275301.